\small

\begin{thebibliography}{99}


\bibitem{Kac0}
V.G. Kac, Simple irreducible graded Lie algebras of finite growth,
{\it Izv. Akad. Nauk SSSR. Ser. Mat.}, {\bf 32}, 1323-1367 (1968).

\bibitem{Moody}
R.V. Moody, A new class of Lie algebras, {\it J. Algebra} {\bf
10}, 211-230 (1968).

\bibitem{Kac}
V.G. Kac, Infinite-dimensional Lie Algebras (Cambridge University
Press, Cambridge, 1990).

\bibitem{FS}
J. Fuchs and C. Schweigert, Symmetries, Lie algebras and
Representations. A graduate course for physicists (Cambridge
University Press, Cambridge, 1997).

\bibitem{Nik}
V.V. Nikulin,  On classification of hyperbolic systems of roots of
rank 3,  Proceedings of the Steklov Institute of Mathematics, {\bf
230} (3), 1–241 (2000) (in Russian).

\bibitem{HPS}
M. Henneaux, D. Persson and P. Spindel, Spacelike Singularities
and Hidden Symmetries of Gravity,  {\it Living Rev. Relativity},
{\bf 11}, 1-228, (2008).

\bibitem{Sac}
C. Saclioglu, Dynkin diagram for hyperbolic Kac-Moody algebras,
{\it J. Phys. A} {\bf 22} (18), 3753-3769 (1989).

\bibitem{BS}
S. de Buyl and C. Schomblond, Hyperbolic Kac Moody algebras and
Einstein biliards, {\it J. Math. Phys.}, {\bf 45} (12) 4464-4492
(2004).

\bibitem{FF}
A. Feingold and I.B. Frenkel, {\it Math. Ann.} {\bf 263}, 87
(1983).

\bibitem{J}
B. Julia, in {\it Lectures in Applied  Mathematics}, {\bf 21}, 355
(1985).

\bibitem{Miz}
S. Mizoguchi, $E_{10}$ Symmetry in One Dimensional  Supergravity,
{\it Nucl. Phys.} {\bf B 528},  238-264 (1998); hep-th/9703160.

\bibitem{Nic1}
H. Nicolai, {\it Phys. Lett. } {\bf B 276}, 333 (1992).

\bibitem{Moore}
G. Moore, String duality, automorphic forms, and generalized
Kac-Moody algebras, {\it Nucl. Phys. Proc. Suppl. } {\bf 67},
56-67 (1998); hep-th/9710198.

\bibitem{DamH3}
T. Damour and M. Henneaux, $E_{10}$, $BE_{10}$ and Arithmetical
Chaos in Superstring Cosmology, {\it Phys. Rev. Lett. } {\bf 86},
4749-4752 (2001); hep-th/0012172.

\bibitem{DHN}
T. Damour, M. Henneaux and H. Nicolai, Cosmological billiards,
topical review, {\it Class. Quantum  Grav.} {\bf 20}, R145-R200
(2003); hep-th/0212256.

\bibitem{IM-brev}
V.D. Ivashchuk and  V.N. Melnikov, On billiard approach in
multidimensional cosmological models, {\it Grav.  Cosmol.} {\bf
15}, No. 1, 49-58 (2009); arXiv: 0811.2786.

\bibitem{W}
P. West, $E11$ and $M$ theory, {\it Class. Quant. Grav.} {\bf 18}
(2001) 4443; hep-th/0104081.

\bibitem{SW}
I. Schnakenburg and P. West, Kac-Moody symmetries of IIB
supergravity, {\it Phys. Lett.} {\bf B 517} (2001) 421,
hep-th/0107181.

\bibitem{IMBl}
V.D. Ivashchuk and V.N. Melnikov, Majumdar-Papapetrou Type
Solutions in Sigma-model and Intersecting p-branes, {\it Class.
Quantum Grav.} {\bf 16}, 849 (1999); hep-th/9802121.

\bibitem{GrI}
M.A. Grebeniuk and V.D. Ivashchuk, Sigma-model solutions and
intersecting p-branes related to Lie algebras, {\it Phys. Lett. }
{\bf B 442}, 125-135 (1998); hep-th/9805113.

\bibitem{IKM}
V.D. Ivashchuk, S.-W. Kim and V.N. Melnikov, Hyperbolic Kac-Moody
algebra from intersecting $p$-branes,  {\it J. Math. Phys.} {\bf
40}, 4072-4083 (1998); {\it Erratum ibid.} {\bf 42}, No 11 (2001);
hep-th/9803006.

\bibitem{IK}
V.D. Ivashchuk and S.-W. Kim, Solutions with  intersecting
p-branes related to Toda chains, {\it J. Math.  Phys.} {\bf 41}
(1), 444-460 (2000); hep-th/9907019.

\bibitem{IMp1}
V.D. Ivashchuk and V.N. Melnikov, P-brane black Holes for General
Intersections,  {\it Grav.  Cosmol.} {\bf 5}, No 4 (20), 313-318
(1999); gr-qc/0002085.

\bibitem{IMp2}
V.D. Ivashchuk and V.N. Melnikov, Black hole p-brane solutions for
general intersection rules, {\it Grav.  Cosmol.} {\bf 6}, No 1
(21), 27-40 (2000); hep-th/9910041.

\bibitem{IMp3}
V.D. Ivashchuk and V.N. Melnikov, Toda  p-brane black holes and
polynomials related to Lie algebras, {\it Class.  Quantum Grav.}
{\bf 17}, 2073-2092 (2000); math-ph/0002048.

\bibitem{IM-top}
V.D. Ivashchuk and V.N. Melnikov,  Exact solutions in
multidimensional gravity with antisymmetric forms, topical review,
{\it Class. Quantum Grav.} {\bf 18},  R82-R157 (2001);
hep-th/0110274.

\bibitem{IMJ}
V.D. Ivashchuk and V.N. Melnikov, Multidimensional classical and
quantum cosmology  with intersecting $p$-branes, {\it J. Math.
Phys.}, {\bf 39}, 2866-2889 (1998); hep-th/9708157.

\bibitem{IMC}
V.D. Ivashchuk and V.N. Melnikov, Sigma-model for the Generalized
Composite p-branes, {\it Class. Quantum Grav.} {\bf 14}, 3001-3029
(1997); Corrigenda {\it ibid.} {\bf 15 } (12), 3941 (1998);
hep-th/9705036.

\bibitem{CJS}
E. Cremmer, B. Julia, and J. Scherk, {\it Phys. Lett.} {\bf B76},
409 (1978).

\bibitem{EH}
F. Englert and L. Houart, $G^{+++}$ invariant formulation of
gravity and M-theories:  Exact intersecting brane solutions, {\it
JHEP} 0405, 059 (2004); hep-th/0405082.

\bibitem{IMb2}
V.D. Ivashchuk and V.N. Melnikov, Billiard representation for
multidimensional cosmology with multicomponent  perfect fluid near
the singularity, {\it Class. Quantum Grav.} {\bf 12}, 809 (1995);
gr-qc/9407028.

\bibitem{IMb3}
V.D. Ivashchuk and V.N. Melnikov, Billiard representation for
multidimensional cosmology with intersecting p-branes near the
singularity, {\it J. Math. Phys.},  {\bf 41}, No 8, 6341-6363
(2000); hep-th/9904077.

\bibitem{AR}
I.Ya. Aref'eva and O.A. Rytchkov, Incidence Matrix Description of
Intersecting p-brane Solutions, {\it Am. Math. Soc. Transl.} {\bf
201}, 19-38 (2000); hep-th/9612236.

\bibitem{AEH}
R. Argurio, F. Englert and L. Hourant, Intersection rules for
$p$-branes,  {\it Phys. Lett. }  {\bf B 398}, 61 (1997);
hep-th/9701042.

\bibitem{GH}
G.W. Gibbons and S.W. Hawking, {\it Phys. Rev.} {\bf D 15}, 2752
(1977).

\bibitem{IMZ}
V.D. Ivashchuk, V.N. Melnikov and A.I. Zhuk, {\it Nuovo Cimento }
{\bf B 104}, 575  (1989).

\bibitem{GOW}
M. Gaberdiel, D. Olive and P. West, A class of Lorentzian
Kac-Moody  algebras, {\it Nucl. Phys.} {\bf B 645} (2002) 403;
hep-th/0205068.

\bibitem{AIV}
I.Ya. Aref'eva, M.G. Ivanov and I.V. Volovich, Non-extremal
intersecting p-branes in various dimensions, {\it  Phys. Lett. }
{\bf B 406}, 44-48 (1997); hep-th/9702079.

\bibitem{Oh}
N. Ohta, Intersection rules for non-extreme p-branes, {\it Phys.
Lett. } {\bf B 403}, 218-224 (1997); hep-th/9702164.

\bibitem{BIM}
K.A. Bronnikov, V.D. Ivashchuk and V.N. Melnikov, The
Reissner-Nordstr\"om Problem for Intersecting  Electric and
Magnetic p-Branes,  {\it Grav. Cosmol.} {\bf 3}, No 3 (11),
203-212 (1997); gr-qc/9710054.

\bibitem{KKLP}
N. Khviengia, Z. Khviengia, H. L\"u and C.N. Pope, Towards a field
theory of F-theory, {\it Class. Quantum Grav.}  {\bf 15}, 759-773
(1998); hep-th/9703012.

\bibitem{LMMP}
H. L\"u, J. Maharana, S. Mukherji  and C.N. Pope, Cosmological
Solutions, p-branes and the Wheeler De Witt Equation, {\it Phys.
Rev.} {\bf D 57 }, 2219-2229 (1997); hep-th/9707182.

\bibitem{Br1}
K.A. Bronnikov, Block-orthogonal Brane systems, Black Holes and
Wormholes,  {\it Grav.  Cosmol.} {\bf 4}, No 1 (13), 49 (1998);
hep-th/9710207.

\bibitem{IMJ2}
V.D. Ivashchuk and V.N. Melnikov, Multidimensional cosmological
and  spherically symmetric solutions with intersecting p-branes.
In  Lecture Notes in Physics, Vol. 537, "Mathematical and Quantum
Aspects of Relativity and Cosmology Proceedings of the Second
Samos Meeting on Cosmology, Geometry and Relativity held at
Pythagoreon, Samos, Greece, 1998, eds: S. Cotsakis, G.W. Gibbons.,
Berlin, Springer, 2000; gr-qc/9901001.

\bibitem{GrIK}
M.A. Grebeniuk, V.D. Ivashchuk and S.-W. Kim, Black-brane
solutions for $C_2$ algebra,  {\it J. Math. Phys.} {\bf 43},
6016-6023 (2002);   hep-th/0111219.

\bibitem{GrIM}
M.A. Grebeniuk, V.D. Ivashchuk and V.N. Melnikov, Black-brane
solution for $A_3$ algebra,  {\it Phys. Lett. }, {\bf B 543},
98-106 (2002);  hep-th/0208083.

\bibitem{IM-SIGMA}
V.D. Ivashchuk, V.N. Melnikov,  On brane solutions related to
non-singular Kac-Moody algebras, {\it SIGMA} {\bf 5}(2009), 070,
34 pages; arxiv: 0810.0196.

\bibitem{I-2010}
V.D. Ivashchuk, On analogues of black brane solutions in the model
with multicomponent anisotropic fluid, {\it Phys. Lett. } {\bf B
693},  399-403 (2010); arxiv: 1001.4053.

\end{thebibliography}

\documentclass{article}
\usepackage{amssymb}
\usepackage{latexsym}

\def\nqq{\hspace{-2em}}

\newcommand{\beq}[1]{\begin{equation}\label{#1}}
\newcommand{\eeq}{\end{equation}}
\newcommand{\bear}[1]{\begin{eqnarray}\label{#1}}
\newcommand{\ear}{\end{eqnarray}}
\newcommand{\nn}{\nonumber}

\catcode`\@=11 \@addtoreset{equation}{section}\catcode`\@=12

\def\barr{\left(\begin{array}}
\def\earr{\end{array}\right)}

\def\mm{\\ \nqq}

\newcommand{\N}{ {\mathbb N} }
\newcommand{\R}{ {\mathbb R} }

\newcommand{\sign}{ \mbox{\rm sign} }

\newcommand{\eps}{ \varepsilon }
\newcommand{\rank}{ \mbox{\rm rank} }

\newcommand{\p}{\partial}

\newcommand{\tri}{\Delta}

\newcommand{\fnm}{\footnotemark}
\newcommand{\fnt}{\footnotetext}

\begin{document}

\begin{center}

\large \bf Black brane solutions related to non-singular Kac-Moody
algebras

\end{center}

\vspace{0.3truecm}

\begin{center}

\normalsize\bf V. D. Ivashchuk\fnm[1]\fnt[1]{e-mail:
ivashchuk@mail.ru} and V. N. Melnikov\fnm[2]\fnt[2]{e-mail:
melnikov@phys.msu.ru},

\vspace{0.3truecm}

\it Center for Gravitation and Fundamental Metrology, VNIIMS,
Ozyornaya ul. 46, Moscow 119361, Russia

\it Institute of Gravitation and Cosmology, Peoples' Friendship
University of Russia,   Miklukho-Maklaya ul. 6, Moscow 117198,
Russia

\end{center}

\begin{abstract}

A multidimensional gravitational model containing scalar fields
and antisymmetric forms is considered. The manifold is chosen in
the form $M = M_0 \times M_1 \times \ldots \times M_n$, where
$M_i$ are Einstein spaces ($i \geq 1$). The sigma-model approach
and exact solutions with intersecting composite branes (e.g.,
solutions with harmonic functions and black brane ones) with
intersection rules related to non-singular Kac-Moody (KM) algebras
(e.g. hyperbolic ones) are considered.  Some examples  of black
brane solutions are  presented, e.g., those corresponding to
hyperbolic KM algebras: $H_2(q,q)$ ($q > 2$), $HA_2^{(1)} =
A_2^{++}$ and to the Lorentzian KM algebra  $P_{10}$.

\end{abstract}

\newpage

\section{\bf Introduction}
\setcounter{equation}{0}

In this paper, we consider certain classes of black brane
solutions related to non-singular   Kac-Moody  algebras. At
present, Kac-Moody (KM) Lie algebras \cite{Kac0,Moody} play
rather an important  role in different areas of mathematical
physics (see \cite{Kac,FS,Nik,HPS} and references therein).  We
recall that KM Lie algebra is a Lie algebra  generated by the
relations \cite{Kac}
 \bear{0.2}
[h_i,h_j] =0, \quad [e_i,f_j] = \delta_{ij} h_j, \mm \label{0.3}
[h_i,e_j] = A_{ij} e_j, \quad [h_i,f_j] = -A_{ij} f_j, \mm
\label{0.4} ({\rm ad} e_i)^{1- A_{ij}}(e_j) = 0 \quad (i \neq j),
\mm \label{6.5} ({\rm ad} f_i)^{1- A_{ij}}(f_j) = 0 \quad (i \neq
j). \ear

Here $A = (A_{ij})$ is a generalized  Cartan matrix,     $i,j = 1,
\ldots, r$, and   $r$ is the rank of the KM algebra. It means that
all  $A_{ii} =   2$;   $A_{ij}$ for $i \neq j$ are non-positive
integers and $A_{ij} =0 $ implies $A_{ji} =0$.    In what follows,
the matrix $A$ is restricted to    be non-degenerate (i.e. $\det A
\neq 0$) and symmetrizable,    i.e. $A = B {\cal D}$, where $B$ is
a symmetric matrix and    ${\cal D}$ is an invertible diagonal
matrix.

If $A$ is positive-definite (the Euclidean case) we get ordinary
finite-dimensional Lie algebras \cite{Kac,FS}. For non-Euclidean
signatures of $A$     all KM algebras are infinite-dimensional.
Among these the Lorentzian     KM algebras  with pseudo-Euclidean
signatures  $(-,+, \ldots, +)$  of the Cartan matrix $A$ are of
current interest  since they contain a subclass  of the so-called
hyperbolic KM algebras widely used in modern mathematical physics.
Hyperbolic KM algebras are by definition Lorentzian Kac-Moody
algebras with the property that, removing any node from their
Dynkin  diagram leaves one with a Dynkin diagram of the affine or
finite type.  The hyperbolic KM algebras were completely
classified in  \cite{Sac,BS}. They have  rank $2 \leq r \leq 10$.
For $r \geq 3$  there is a finite number of hyperbolic algebras.
For rank 10, there  are four algebras, known as $E_{10}$,
$BE_{10}$, $CE_{10}$  and $DE_{10}$. Hyperbolic KM algebras
appeared in ordinary gravity  \cite{FF} (${\cal F}_3 = AE_3 =
H_3$),  supergravity: \cite{J,Miz} ($E_{10}$),  \cite{Nic1}
(${\cal F}_3$), strings \cite{Moore}, oscillating behaviour near
the singularity \cite{DamH3} (see also \cite{DHN,IM-brev}), etc.

It has been  proposed by P. West that the Lorentzian
(non-hyperbolic) KM algebra $E_{11}$   is responsible for  a
hidden algebraic structure   characterizing  $11D$ supergravity
\cite{W}.  The same very extended algebra occurs in $IIA$ \cite{W}
and $IIB$ supergravities \cite{SW}.

Here we briefly consider another possibility    of utilizing
non-singular (e.g. hyperbolic) KM algebras,  suggested  in our
three papers \cite{IMBl,GrI,IKM}.    This possibility also
implicitly assumed in    \cite{IK,IMp1,IMp2,IMp3,IM-top} is
related to certain classes of  exact solutions describing
intersecting composite branes  in a multidimensional gravitational
model containing scalar fields  and antisymmetric forms  defined
on (warped) product  manifolds $M = M_0 \times M_1 \times \ldots
\times M_n$, where  $M_i$ are Ricci-flat spaces ($i \geq 1$). From
a pure mathematical point of view, these solutions  may be
obtained from sigma-models or Toda chains  corresponding to
certain non-singular KM algebras.  The information about a
(hidden) KM algebra is encoded in intersection rules which relate
the dimensions of brane intersections  with non-diagonal
components of the generalized Cartan matrix $A$ \cite{IMJ}. We
deal here with generalized Cartan matrices of the form
\beq{1.1}
A_{ss'} \equiv \frac{2(U^s,U^{s'})}{(U^{s'},U^{s'})},
\eeq
 $s,s'\in S$, with $(U^s,U^s)\neq 0$, for all $s\in S$ ($S$ is a
finite set). Here $U^s$ are the so-called brane (co-)vectors. They
are linear functions on $\R^N$, where $N = n+l$ and  $l$ is the
number of scalar fields. The indefinite   scalar product $(.,.)$
\cite{IMC} is defined on $(\R^N)^{*}$ and has the signature
$(-1,+1, \ldots, +1)$  if all scalar fields have positive kinetic
terms, i.e. there  are no phantoms (or ghosts).   The matrix   $A$
is symmetrizable.  $U^s$-vectors may be put in one-to-one
correspondence with simple roots  $\alpha_s$ of the generalized KM
algebra after a suitable normalization.

For   $D=11$ supergravity  \cite{CJS} and ten-dimensional  $IIA$,
$IIB$ supergravities all $(U^s,U^s) = 2$ \cite{IMJ,EH},  and
corresponding KM algebras are simply laced. It was shown in our
papers \cite{IMb2,IMb3} that the inequality $(U^s,U^s)> 0$ is a
necessary condition for the formation  of a billiard wall (if one
approaches the singularity) by the $s$-th  matter source (e.g., a
fluid component or a brane).

The scalar products  for brane vectors $U^s$  were found in
\cite{IMC}
\beq{1.2} (U^s,U^{s'})= d_{ss'}+\frac{d_s d_{s'}}{2-D}+
\chi_s \chi_{s'} <\lambda_{s}, \lambda_{s'}> , \eeq
where $d_s$
and $d_{s'}$ are the dimensions of  brane world volumes
corresponding to branes $s$ and $s'$, respectively, $d_{ss'}$ is
the dimension of the intersection of  brane world volumes, $D$ is
the total space-time dimension, $\chi_s = + 1, -1$ for electric or
magnetic brane respectively, and $<\lambda_{s}, \lambda_{s'}>$ is
the non-degenerate scalar product of the $l$-dimensional dilatonic
coupling vectors $\lambda_{s}$ and $\lambda_{s'}$  corresponding
to branes $s$ and $s'$.

The relations (\ref{1.1}), (\ref{1.2}) determine the brane
intersection rules \cite{IMJ}
\beq{1.3} d_{s s'}= d_{s s'}^{o} +
\frac12 K_{s'} A_{s s'},
\eeq
 $s \ne s'$, where $K_s = (U^s,U^s)$ and
\beq{1.4} d_{s s'}^{o} =
\frac{d_s d_{s'}}{D-2} -  \chi_s\chi_{s'} <\lambda_{s},
\lambda_{s'}>
\eeq
 is the dimension of the so-called orthogonal (or ($A_1 \oplus
A_1$)-) intersection of branes following from the orthogonality
condition  \cite{IMC}
\beq{1.5}
(U^s,U^{s'})= 0,
\eeq
 $s \ne s'$.

The relations (\ref{1.2}) and (\ref{1.4}) were derived in
\cite{IMC} under rather general assumptions: the branes were
composite, the number of scalar fields $l$ was arbitrary as well
as the signature of the bilinear form $<.,.>$ (or, equivalently,
the signature of the kinetic term for scalar fields),  Ricci-flat
factor spaces $M_i$ had arbitrary dimensions $d_i$ and signatures.
The intersection rules  (\ref{1.4}) appeared earlier in
\cite{AR,AEH} for all $d_i =1$ ($i > 0$) and $<.,.>$ being
positive-definite.

\section{\bf The model}
\setcounter{equation}{0}

\subsection{The action }

We consider a model governed by the action
\bear{2.1} S =&&
\frac{1}{2\kappa^{2}} \int_{M} d^{D}z \sqrt{|g|} \{ {R}[g] - 2
\Lambda - h_{\alpha\beta} g^{MN} \partial_{M} \varphi^\alpha
\partial_{N} \varphi^\beta
 \\ \nn
&& - \sum_{a \in {\bf \Delta}} \frac{\theta_a}{n_a!} \exp[ 2
\lambda_{a} (\varphi) ] (F^a)^2_g \}  + S_{GH}, \ear
 where $g = g_{MN} dz^{M} \otimes dz^{N}$ is the metric on the
manifold $M$, ${\dim M} = D$, $\varphi=(\varphi^\alpha)\in \R^l$
is a vector of dilatonic scalar fields, $(h_{\alpha \beta})$ is a
non-degenerate symmetric  $l\times l$ matrix ($l\in \N$),
$\theta_a  \neq 0$,
$$ F^a =  dA^a =  \frac{1}{n_a!} F^a_{M_1
\ldots M_{n_a}} dz^{M_1} \wedge \ldots \wedge dz^{M_{n_a}}$$
 is an $n_a$-form ($n_a \geq 2$) on the $D$-dimensional manifold $M$,
$\Lambda$ is a cosmological constant and $\lambda_{a}$ is a
$1$-form on $\R^l$: $\lambda_{a} (\varphi) =\lambda_{a
\alpha}\varphi^\alpha$, $a \in {\bf \Delta}$, $\alpha=1,\ldots,l$.
In (\ref{2.1}) we denote $|g| = |\det (g_{MN})|$, $(F^a)^2_g =
F^a_{M_1 \ldots M_{n_a}} F^a_{N_1 \ldots N_{n_a}} g^{M_1 N_1}
\ldots g^{M_{n_a} N_{n_a}}$, $a \in {\bf \Delta}$, where ${\bf
\Delta}$ is  some finite set (for example, of positive integers),
and $S_{\rm GH}$ is the standard Gibbons-Hawking boundary term
\cite{GH}. In models with one time all $\theta_a  =  1$  and the
signature of the metric is $(-1,+1, \ldots, +1)$;  $\kappa^{2}$ is
the multidimensional gravitational constant.

\subsection{The Ansatz for composite  branes }

Consider the manifold
\beq{2.10} M = M_{0}  \times M_{1} \times
\ldots \times M_{n}, \eeq
 with the metric
\beq{2.11} g= e^{2{\gamma}(x)} \hat{g}^0  + \sum_{i=1}^{n}
e^{2\phi^i(x)} \hat{g}^i , \eeq
 where $g^0  = g^0 _{\mu \nu}(x)
dx^{\mu} \otimes dx^{\nu}$ is an arbitrary metric with any
signature on the manifold $M_{0}$ and $g^i  = g^{i}_{m_{i}
n_{i}}(y_i) dy_i^{m_{i}} \otimes dy_i^{n_{i}}$ is a metric on
$M_{i}$  satisfying the equation
\beq{2.13}
 R_{m_{i}n_{i}}[g^i ] = \xi_{i} g^i_{m_{i}n_{i}},
\eeq
 $m_{i},n_{i}=1, \ldots, d_{i}$; $\xi_{i}= {\rm const}$,
$i=1,\ldots,n$. Here $\hat{g}^{i} = p_{i}^{*} g^{i}$ is the
pullback of the metric $g^{i}$  to the manifold  $M$ by the
canonical projection: $p_{i} : M \rightarrow  M_{i}$, $i =
0,\ldots, n$. Thus, $(M_i, g^i )$  are Einstein spaces, $i =
1,\ldots, n$. The functions $\gamma, \phi^{i} : M_0 \rightarrow \R
$ are smooth. We denote $d_{\nu} = {\rm dim} M_{\nu}$; $\nu = 0,
\ldots, n$;  $D = \sum_{\nu = 0}^{n} d_{\nu}$. We assume all
manifolds $M_{\nu}$, $\nu = 0,\ldots, n$, to be oriented and
connected. Then the volume $d_i$-form
\beq{2.14} \tau_i  \equiv
\sqrt{|g^i(y_i)|} \ dy_i^{1} \wedge \ldots \wedge dy_i^{d_i}, \eeq
 and signature parameter
\beq{2.15} \varepsilon(i)  \equiv {\rm
sign}( \det (g^i_{m_i n_i})) = \pm 1
\eeq
 are correctly defined for all $i=1,\ldots,n$.

Let $\Omega = \Omega(n)$  be a set of all non-empty subsets of $\{
1, \ldots,n \}$. The number of elements in $\Omega$ is $|\Omega| =
2^n - 1$. For any $I = \{ i_1, \ldots, i_k \} \in \Omega$, $i_1 <
\ldots < i_k$, we denote
\bear{2.16}
\tau(I) \equiv \hat{\tau}_{i_1}  \wedge \ldots \wedge \hat{\tau}_{i_k},  \\
\label{2.17}
\eps(I) \equiv \eps(i_1) \ldots \eps(i_k),  \\
\label{2.18}
M_{I} \equiv M_{i_1}  \times  \ldots \times M_{i_k}, \\
\label{2.19} d(I) \equiv  \sum_{i \in I} d_i. \ear

Here $\hat{\tau}_{i} = p_{i}^{*} \hat{\tau}_{i}$ is the pullback
of the form $\tau_i$  to the manifold  $M$ by the canonical
projection: $p_{i} : M \rightarrow  M_{i}$, $i = 1,\ldots, n$. We
also put $\tau(\emptyset)= \eps(\emptyset)= 1$ and
$d(\emptyset)=0$.

For fields of forms we consider the following composite
electromagnetic ansatz:
 \beq{2.1.1}
F^a=\sum_{I\in\Omega_{a,e}}{\cal F}^{(a,e,I)}+
\sum_{J\in\Omega_{a,m}}{\cal F}^{(a,m,J)}
\eeq
 where
\bear{2.1.2}
{\cal F}^{(a,e,I)}=d\Phi^{(a,e,I)}\wedge\tau(I), \\
\label{2.1.3} {\cal F}^{(a,m,J)}=
e^{-2\lambda_a(\varphi)}*(d\Phi^{(a,m,J)} \wedge\tau(J))
\ear
 are  elementary forms of electric and magnetic types, respectively,
$a \in {\bf \Delta}$, $I \in \Omega_{a,e}$, $J \in \Omega_{a,m}$
and  $\Omega_{a,v} \subset \Omega$, $v = e,m$. In (\ref{2.1.3})
$*=*[g]$ is the Hodge operator on $(M,g)$.

For scalar functions we put
\beq{2.1.5}
\varphi^\alpha=\varphi^\alpha(x), \quad \Phi^s=\Phi^s(x),
\eeq
 $s\in S$. Thus $\varphi^{\alpha}$ and $\Phi^s$ are functions on
$M_0$.

Here and below
\beq{2.1.6}
S=S_e \sqcup S_m, \quad S_v= \sqcup_{a
\in {\bf \Delta}} \{ a \} \times \{v \} \times \Omega_{a,v},
\eeq
 $v=e,m$. Here and in what follows $\sqcup$ means the union of
non-intersecting sets. The set $S$ consists of elements
$s=(a_s,v_s,I_s)$, where $a_s \in {\bf \Delta}$ is color index,
$v_s = e, m$ is the electro-magnetic index and the set $I_s \in
\Omega_{a_s,v_s}$ describes the location of a brane.

Due to (\ref{2.1.2}) and (\ref{2.1.3})
\beq{2.1.7}
d(I)=n_a-1,
\quad d(J)=D-n_a-1,
\eeq
 for  $I \in \Omega_{a,e}$ and $J \in \Omega_{a,m}$ (i.e., in the electric and
magnetic case, respectively).

\subsection{The sigma model}

Let $d_0 \neq 2$ and
\beq{2.2.1} \gamma=\gamma_0(\phi) \equiv
\frac1{2-d_0}\sum_{j=1}^nd_j\phi^j,
\eeq
 i.e., the generalized harmonic gauge (frame) is used.

We put two restrictions on the  sets of branes that guarantee the
block-diagonal form of the  energy-momentum tensor and the
existence of the sigma-model representation (without additional
constraints):

\beq{2.2.2a}
{\bf (R1)} \quad d(I \cap J) \leq d(I) - 2,
\eeq
 for any $I,J \in\Omega_{a,v}$, $a \in {\bf \Delta}$, $v= e,m$ (here
$d(I) = d(J)$) and

\beq{2.2.3a} {\bf (R2)} \quad d(I \cap J) \neq 0 \ for \ d_0 = 1,
\qquad d(I \cap J) \neq 1 \quad for \ d_0  = 3. \eeq

It was proved in \cite{IMC} that equations of motion for the model
(\ref{2.1}) and the Bianchi identities:

\beq{2.2.6} d{\cal F}^s=0,
\eeq
 $s \in S_m$, for fields from (\ref{2.11}),
(\ref{2.1.1})-(\ref{2.1.5}), under the restrictions (\ref{2.2.2a})
and (\ref{2.2.3a}), are equivalent to the equations of motion of
the $\sigma$-model governed by the action

\bear{2.2.7}
S_{\sigma 0} = && \frac{1}{2 \kappa_0^2} \int
d^{d_0}x\sqrt{|g^0|}\biggl\{R[g^0]-\hat G_{AB}
g^{0\mu\nu}\p_\mu\sigma^A\p_\nu\sigma^B  \\ \nn &&-\sum_{s\in
S}\eps_s \exp{(-2U_A^s\sigma^A)} g^{0\mu\nu}
\p_\mu\Phi^s\p_\nu\Phi^s - 2V \biggr\},
\ear
 where $(\sigma^A)=(\phi^i,\varphi^\alpha)$, $k_0 \neq 0$, the
index set  $S$ is defined in (\ref{2.1.6}),

\beq{2.2.8}
V = {V}(\phi)  = \Lambda e^{2 {\gamma_0}(\phi)}
-\frac{1}{2} \sum_{i =1}^{n} \xi_i d_i e^{-2 \phi^i + 2
{\gamma_0}(\phi)}
\eeq
 is the potential,
\beq{2.2.9}
(\hat G_{AB})= \barr{cc}
G_{ij}& 0\\
0& h_{\alpha\beta}
\earr \eeq
 is the target space metric with
 \beq{2.2.10}
 G_{ij}= d_i \delta_{ij}+\frac{d_i d_j}{d_0-2}
\eeq
and co-vectors
\beq{2.2.11}
U_A^s =   U_A^s \sigma^A =
\sum_{i \in I_s} d_i \phi^i - \chi_s \lambda_{a_s}(\varphi), \quad
(U_A^s) = (d_i \delta_{iI_s}, -\chi_s \lambda_{a_s \alpha}),
\eeq
 $s=(a_s,v_s,I_s)$. Here $\chi_e=+1$ and $\chi_m=-1$;

\beq{2.2.12}
\delta_{iI}=\sum_{j\in I}\delta_{ij}
\eeq
 is an indicator of $i$ belonging to $I$: $\delta_{iI}=1$ for $i\in
I$ and $\delta_{iI}=0$ otherwise; and

\bear{2.2.13}
\eps_s=(-\eps[g])^{(1-\chi_s)/2}\eps(I_s)
\theta_{a_s},
\ear
 $s\in S$, $\eps[g]\equiv\sign\det(g_{MN})$. More explicitly,
(\ref{2.2.13}) reads

\beq{2.2.13a}
\eps_s=\eps(I_s) \theta_{a_s} \ {\rm for} \ v_s = e;
\qquad \eps_s = -\eps[g] \eps(I_s) \theta_{a_s},
 \ {\rm for} \ v_s  = m.
\eeq

For finite internal space volumes $V_i$ (e.g. compact $M_i$) and
electric $p$-branes  (i.e. all $\Omega_{a,m} = \emptyset$) the
action (\ref{2.2.7}) coincides with the action (\ref{2.1}) when
$\kappa^{2} = \kappa^{2}_0 \prod_{i=1}^{n} V_i$.

\section{Solutions governed by harmonic functions}

\subsection{Solutions with a block-orthogonal
             set of $U^s$ and Ricci-flat factor-spaces}

Here we consider a special class of solutions to the equations of
motion governed by several harmonic functions, where all factor
spaces are Ricci-flat and the cosmological constant is zero, i.e.,
$\xi_i = \Lambda = 0$, $i = 1,\ldots,n$. In certain situations
these solutions describe extremal black branes charged by fields
of forms.

The solutions crucially depend upon  scalar products of
$U^s$-vectors $(U^s,U^{s'})$; $s,s' \in S$, where
\beq{3.1.1}
(U,U')=\hat G^{AB} U_A U'_B, \eeq for $U = (U_A), U' = (U'_A) \in
\R^N$, $N = n + l$ and \beq{3.1.2} (\hat G^{AB})
=\left(\begin{array}{cc}
 G^{ij}& 0\\
 0     & h^{\alpha\beta}
\end{array}\right)
\eeq
 is the inverse matrix  to  the matrix (\ref{2.2.9}). As
in \cite{IMZ}, we have
\beq{3.1.3}
G^{ij}=\frac{\delta^{ij}}{d_i}+\frac1{2-D},
\eeq
 $i,j=1,\dots,n$.

The scalar products (\ref{3.1.1}) of the vectors $U^s$  were
calculated in \cite{IMC} and are given by
\beq{3.1.4}
(U^s,U^{s'})=d(I_s\cap I_{s'})+\frac{d(I_s)d(I_{s'})}{2-D}+
\chi_s\chi_{s'}\lambda_{a_s \alpha} \lambda_{a_{s'} \beta}
h^{\alpha \beta}, \eeq
where
$(h^{\alpha\beta})=(h_{\alpha\beta})^{-1}$, and $s=(a_s,v_s,I_s)$,
$s'=(a_{s'},v_{s'},I_{s'})$ belong to $S$. This relation is  very
important one since it encodes  brane data (e.g., intersections)
via the  scalar products of $U$-vectors.

Let
\beq{3.1.5} S=S_1 \sqcup \dots \sqcup S_k,
\eeq
$S_i\ne\emptyset$, $i=1,\dots,k$, and
\beq{3.1.6} (U^s,U^{s'})=0
\eeq
for all $s\in S_i$, $s'\in S_j$, $i\ne j$; $i,j=1,\dots,k$.
The relation (\ref{3.1.5}) means that the set $S$ is a union of
$k$ non-intersecting (non-empty) subsets $S_1,\dots,S_k$.
According to (\ref{3.1.6}) the set of vectors $(U^s, s \in S)$ has
a block-orthogonal structure with respect to the scalar product
(\ref{3.1.1}), i.e., it  splits into $k$ mutually orthogonal
blocks $(U^s, s \in S_i)$, $i=1,\dots,k$.

Here we consider exact solutions in the model (\ref{2.1}) where
vectors $(U^s,s\in S)$ obey the block-orthogonal decomposition
(\ref{3.1.5}), (\ref{3.1.6}) with scalar products defined in
(\ref{3.1.4}) \cite{IMBl}. These solutions were obtained from the
corresponding solutions to the $\sigma$-model equations of motion
\cite{IMBl}.

{\bf Proposition  1.} {\em Let $(M_0,g^0)$ be Ricci-flat:
 $R_{\mu\nu}[g^0]=0$. Then the field configuration

\beq{3.1.7}
 g^0, \qquad \sigma^A=\sum_{s\in
S}\eps_sU^{sA}\nu_s^2\ln H_s, \qquad \Phi^s=\frac{\nu_s}{H_s},
\eeq
 $s\in S$,
 satisfies the field equations corresponding to the
action (\ref{2.2.7}) with $V=0$ if the real numbers $\nu_s$ obey
the relations

\beq{3.1.8}
\sum_{s'\in S}(U^s,U^{s'})\eps_{s'}\nu_{s'}^2=-1
\eeq
 $s\in S$, the functions $H_s >0$ are harmonic, i.e.
$\tri[g^0]H_s=0$, $s\in S$, and $H_s$  coincide inside the blocks:
$H_s=H_{s'}$ for $s,s'\in S_i$, $i=1,\dots,k$.}

Using the sigma-model solution from Proposition 1 and the
relations for contra-variant components \cite{IMC}:

\beq{2.2.38} U^{si}=\delta_{iI_s}-\frac{d(I_s)}{D-2}, \quad
 U^{s\alpha}=-\chi_s\lambda_{a_s}^\alpha,
\eeq
 $s=(a_s,v_s,I_s)$,
 we get \cite{IMBl}:
\bear{3.1.11} g= \left(\prod_{s \in S}
H_s^{2d(I_s)\eps_s\nu_s^2}\right)^{1/(2-D)} \left\{\hat{g}^0+
\sum_{i=1}^n \left(\prod_{s \in S}
H_s^{2\eps_s\nu_s^2\delta_{iI_s}}\right)
\hat{g}^i\right\}, \\
\label{3.1.14}
\varphi^\alpha= -\sum_{s \in S}\lambda_{a_s}^\alpha
\chi_s
\eps_s\nu_s^2\ln H_s, \\
\label{3.1.15}
F^a=\sum_{s\in S}{\cal F}^s\delta_{a_s}^a,
\ear
 where $i=1,\dots,n$, $\alpha=1,\dots,l$, $a \in {\bf \Delta}$ and

\bear{3.1.16}
{\cal F}^s=\nu_s dH_s^{-1}\wedge\tau(I_s), \mbox{
for } v_s=e, \mm \label{3.1.17} {\cal F}^s=\nu_s
(*_0dH_s)\wedge\tau(\bar I_s), \mbox{ for } v_s=m,
\ear
 $H_s$ are harmonic functions on $(M_0,g^0)$ which coincide inside
the blocks (i.e., $H_s=H_{s'}$ for $s,s'\in S_i$, $i=1,\dots,k$)
and the relations  (\ref{3.1.8}) on the parameters $\nu_s$ are
imposed. The matrix $((U^s,U^{s'}))$ and the parameters $\eps_s$,
$s\in S$, are defined in (\ref{3.1.4}) and (\ref{2.2.13}),
respectively; $\lambda_a^\alpha=  h^{\alpha\beta}\lambda_{a
\beta}$, $*_0=*[g^0]$ is the Hodge operator  on $(M_0,g^0)$ and

\beq{2.2.5I} \bar I = \{1, \dots, n \} \setminus I \eeq is the
dual set.  (In (\ref{3.1.17}) we have redefined the sign of the
parameter $\nu_{s}$.)

\subsection{Solutions related to non-singular KM algebras}

Now we will study the  solutions (\ref{3.1.11})-(\ref{3.1.17}) in
more detail and show that some of them may be related to
non-singular KM algebras. We put
\beq{3.1.2.1}
  K_s \equiv (U^s,U^s)\neq 0
\eeq
 for all $s\in S$ and introduce the  matrix $A=(A_{ss'})$:
\beq{3.1.2.2}
 A_{ss'} \equiv \frac{2(U^s,U^{s'})}{(U^{s'},U^{s'})},
 \eeq
$s,s'\in S$. Here  some ordering in $S$ is assumed.

Using this definition and (\ref{3.1.4}) we obtain the intersection
rules \cite{IMJ}

\beq{3.1.2.3} d(I_{s}\cap I_{s'})=\Delta(s,s')+\frac12 K_{s'} A_{s
s'}, \eeq where $s \ne s'$, and \beq{3.1.D} \Delta(s,s') =
\frac{d(I_s)d(I_{s'})}{D-2} - \chi_s\chi_{s'}\lambda_{a_s \alpha}
\lambda_{a_{s'} \beta} h^{\alpha \beta}
\eeq
 defines the so-called ``orthogonal'' intersection rules \cite{IMC}
(see also  \cite{AR,AEH} for $d_i = 1$).

In $D = 11$ and $D = 10$ ($IIA$ and $IIB$ ) supergravity models,
all $K_s = 2$ and hence the intersection rules (\ref{3.1.2.3}) in
this case have a simpler form \cite{IMJ}:
\beq{3.1.2.3.sl}
d(I_{s}\cap I_{s'})= \Delta(s,s')+ A_{s s'}, \eeq
 where $s \ne s'$, implying $A_{s s'} = A_{s' s}$. The
corresponding KM algebra is simply-laced in this case.

For $\det A \neq 0$ relation (\ref{3.1.8}) may be rewritten in the
equivalent form
\beq{3.1.2.5} - \eps_s\nu_s^2(U^s,U^s)= 2
\sum_{s'\in S} A^{ss'} \equiv b_s,
\eeq
 where $s\in S$, and $(A^{ss'})=A^{-1}$. Thus eq. (\ref{3.1.8})
may be resolved in terms of $\nu_s$ for certain $\eps_s=\pm1$,
$s\in S$. We note that due to $(\ref{3.1.6})$ the matrix $A$ has a
block-diagonal structure  and, hence, for any $i$-th block the set
of parameters $(\nu_s,  s \in S_i)$ depends on the matrix inverse
to the matrix  $(A_{s s'};  s,  s' \in S_i)$.

Now  we consider the one-block case such that the brane
intersections are related to some  non-singular KM algebras.

{\bf  Finite-dimensional Lie algebras \cite{GrI}}

Let $A$ be the Cartan matrix of a simple finite-dimensional Lie
algebra. In this case $A_{ss'}\in\{0,-1,-2,-3\}$, $s\ne s'$. The
elements of the inverse matrix $A^{-1}$ are positive (see Ch. 7 in
\cite{FS}) and hence we get from (\ref{3.1.2.5}) the same
signature relation  as in the  orthogonal case \cite{IMC}:
\beq{3.1.2.20b} \eps_s(U^s,U^s) < 0, \eeq
 $s\in S$.

If all $(U^s,U^{s}) > 0$, we get  $\eps_{s} < 0$,  $s \in S$.
Moreover, all $b_s$ are positive integers: \beq{3.1.2.6}
 b_s = n_s \in \N,
\eeq
 $s \in S$.
The integers $n_s$ coincide with the components of the twice dual
Weyl vector in the basis of simple co-roots (see Ch. 3.1.7 in
(\cite{FS}).

{\bf  Hyperbolic KM algebras}

Let $A$ be a generalized Cartan matrix corresponding to a simple
hyperbolic KM algebra. For   hyperbolic algebras, the following
relations are satisfied
\beq{3.1.2.20}
\eps_s(U^s,U^s) >0, \eeq
 since all $b_s$ are negative in the hyperbolic case \cite{GOW}:
\beq{3.1.2.6b}
b_s  < 0,
\eeq
 where $s\in S$.

For $(U^s,U^{s}) > 0$ we get  $\eps_{s} > 0$, $s \in S$. If
$\theta_{a_s} > 0$ for all $s \in S$, then
\beq{3.1.28a}
\eps(I_s)
= 1   \ {\rm for} \ v_s = e; \qquad \eps(I_s) = - \eps[g] \ {\rm
for} \ v_s = m.
\eeq

For a pseudo-Euclidean metric $g$ all $\eps(I_s) = 1$, and hence
all branes are Euclidean or should contain even number of time
directions: $2,4, \ldots$. For $\eps[g] = 1$ only magnetic branes
may be pseudo-Euclidean.

{\bf $B_D$-models.} The $B_D$-model has the  action \cite{IMJ}
\beq{3.1.2.09}
 S_D=\int
 d^Dz\sqrt{|g|}\biggl\{R[g]+ g^{MN}\p_M\vec\varphi\p_N\vec\varphi-
\sum_{a=4}^{D-7}\frac1{a!}\exp[2\vec\lambda_a\vec\varphi](F^a)^2\biggr\},
\eeq
 where   $\vec\varphi=(\varphi^1,\dots,\varphi^l)\in\R^l$,
$\vec\lambda_a= (\lambda_{a1},\dots,\lambda_{al})\in\R^l$,
$l=D-11$, $\rank F^a=a$, $a=4,\dots,D-7$. Here vectors
$\vec\lambda_a$ satisfy the relations

\bear{3.1.2.010}
\vec\lambda_a\vec\lambda_b=N(a,b)-\frac{(a-1)(b-1)}{D-2} =
\Lambda_{ab}, \\
\label{3.1.2.011} N(a,b)=\min(a,b)-3, \ear $a,b=4,\dots,D-7$ and
$\vec\lambda_{D-7}=-2\vec\lambda_4$. For $D>11$ vectors
$\vec\lambda_4,\dots,\vec\lambda_{D-8}$ are linearly independent.
(It may be verified that the matrix $(\Lambda_{ab})$ is
positive-definite  and hence the set of vectors obeying
(\ref{3.1.2.010}) does exist.)

The model (\ref{3.1.2.09}) contains $l$ scalar fields with a
negative kinetic term (i.e.,
$h_{\alpha\beta}=-\delta_{\alpha\beta}$ in (\ref{2.1})) coupled to
several forms (the number of forms is $(l+1)$) .

For the brane worldvolumes we have the following dimensions (see
(\ref{2.1.7})):   $d(I)=3,\dots,D-8$ for $I\in\Omega_{a,e}$
 and $d(I)=D-5,\dots,6$ for $I \in\Omega_{a,m}$.
Thus there are $(l+1)$ electric and $(l+1)$ magnetic $p$-branes,
$p=d(I)-1$. In $B_D$-model all $K_s = 2$. (For $B_{12}$-model see
\cite{KKLP}.)

{\bf Example 1: $H_2(q_1,q_2)$ algebra \cite{IMBl}.} Let

\beq{3.13}
 A =\barr{cc}
 2& -q_1\\
 -q_2& 2 \earr, \quad q_1q_2>4,
\eeq
 $q_1,q_2\in \N$. This is the Cartan matrix of the hyperbolic KM
algebra $H_{2}(q_1,q_2)$ \cite{Kac}.  From (\ref{3.1.2.5}) we get
\beq{3.14} \eps_s\nu_s^2(U^s,U^s)(q_1q_2-4)=2q_s+4, \eeq
 $s\in\{1,2\}=S$.

An example of the $H_{2}(q,q)$-solution for $B_D$-model  with two
electric $p$-branes  ($p=d_1,d_2$), corresponding to $F^a$ and
$F^b$ fields and intersecting on a time manifold, is as follows
\cite{IMBl}: \bear{3.1.2.33}
g=H^{-2/(q-2)}\hat{g}^0-H^{2/(q-2)}dt\otimes
dt+\hat{g}^1+\hat{g}^2, \mm \label{3.1.2.34a}
F^a=\nu_{1}dH^{-1}\wedge dt\wedge \hat{\tau}_1, \mm
\label{3.1.2.34b} F^b = \nu_{2}dH^{-1}\wedge dt\wedge
\hat{\tau}_2, \mm \label{3.1.2.35}
\vec\varphi=-(\vec\lambda_a+\vec\lambda_b)(q -2)^{-1}\ln H \ear
 where $d_0=3$, $d_1=a - 2$, $a=q+4$, $a < b$, $d_2=b-2$, $D=a+b$,
and $\nu_{1}^2 = \nu_{2}^2 = (q - 2)^{-1}$. The signature
restrictions are : $\eps_1= \eps_2 = -1$. Thus the space-time
$(M,g)$ should contain at least three time directions. The minimal
$D$ is 15. For $D=15$ we get $a =7$, $b =8$, $d_1 =5$, $d_2= 6$
and $q=3$.

\section{Black brane solutions}

In this section we consider spherically symmetric solutions with
$d_0 =1$ and  $M_1 = S^{d_1}$, $g^1 = d \Omega^2_{d_1}$, where $d
\Omega^2_{d_1}$ is the canonical metric on a unit sphere
$S^{d_1}$, $d_1 \geq 2$. The manifold $M_0$ corresponds to a
radial variable $R$. We put also  $M_2 = \R$, $g^2 = - dt \otimes
dt$, i.e.,  $M_2$ is a time manifold and
\beq{5.2.18} 2 \in I_s,
\quad \forall s \in S,
\eeq
 i.e., all branes have a common time direction $t$.

In \cite{IMp1,IMp2,IMp3}, the following solutions with a horizon
were obtained:

\bear{5.2.30}
g= \Bigl(\prod_{s \in S}
 H_s^{2 h_s d(I_s)/(D-2)} \Bigr) \biggl\{ \left(1 -
\frac{2\mu}{R^{\bar{d}}}\right)^{-1} dR \otimes dR + R^2  d
\Omega^2_{d_1}  \\ \nn -  \Bigl(\prod_{s \in S} H_s^{-2 h_s}
\Bigr) \left(1 - \frac{2\mu}{R^{\bar{d}}}\right) dt \otimes dt +
\sum_{i = 3}^{n} \Bigl(\prod_{s \in S} H_s^{-2 h_s \delta_{iI_s}}
\Bigr) \hat{g}^i  \biggr\},
\\  \label{5.2.31}
\exp(\varphi^\alpha)= \prod_{s\in S} H_s^{h_s \chi_s
\lambda_{a_s}^\alpha}, \ear
 where $F^a= \sum_{s \in S} \delta^a_{a_s} {\cal F}^{s}$, and
\beq{5.2.32}
{\cal F}^s= - \frac{Q_s}{R^{d_1}} \left( \prod_{s'
\in S}  H_{s'}^{- A_{s s'}} \right) dR \wedge\tau(I_s), \eeq $s\in
S_e$,

\beq{5.2.33}
{\cal F}^s= Q_s \tau(\bar I_s),
\eeq $s \in S_m$.

Here $Q_s \neq 0$, $h_s =K_s^{-1}$, $s \in S$, and the generalized
Cartan matrix $(A_{s s'})$ is non-degenerate.

The functions $H_s > 0$ obey the equations
\beq{5.3.1}
\frac{d}{dz} \left( \frac{(1 - 2\mu z)}{H_s} \frac{d}{dz} H_s
\right) = \bar B_s \prod_{s' \in S}  H_{s'}^{- A_{s s'}}, \eeq

\bear{5.3.2a}
H_{s}((2\mu)^{-1} -0) = H_{s0} \in (0, + \infty), \\
\label{5.3.2b} H_{s}(+ 0) = 1, \ear
 $s \in S$, where $H_s(z) > 0$, $\mu > 0$, $z = R^{-\bar d} \in (0,
(2\mu)^{-1})$, $\bar d = d_1 -1$ and $\bar B_s = \eps_s K_s Q_s^2/
\bar d^2 \neq 0$.

There exist solutions to eqs. (\ref{5.3.1})-(\ref{5.3.2a}) of
polynomial type. The simplest example occurs in orthogonal case
\cite{BIM,IMJ} (for $d_i = 1$  see also \cite{AIV,Oh}):
$(U^s,U^{s'})= 0$, for  $s \neq s'$, $s, s' \in S$. In this case
$(A_{s s'}) = {\rm diag}(2,\ldots,2)$ is a Cartan matrix of the
semisimple Lie algebra ${\bf A_1} \oplus  \ldots  \oplus  {\bf
A_1}$ and

\beq{5.3.5}
H_{s}(z) = 1 + P_s z
\eeq
 with $P_s \neq  0$, satisfying
\beq{5.3.5a}
P_s(P_s + 2\mu) = -\bar B_s,
\eeq
 $s \in S$.

In \cite{Br1,IMJ2} this solution was generalized to the
block-orthogonal case  (\ref{3.1.5}), (\ref{3.1.6}). In this case
(\ref{5.3.5}) is modified as follows
 \beq{5.3.8} H_{s}(z) = (1 + P_s z)^{b_s}, \eeq
 where $b_s$ are defined as follows
\beq{5.2.20}
 b_s = 2 \sum_{s'
\in S} A^{s s'}
\eeq
 and parameters $P_s$ coincide within blocks, i.e., $P_s = P_{s'}$
for $s, s' \in S_i$, $i =1,\dots,k$. The parameters $P_s \neq 0 $
satisfy the relations \cite{IMJ2,IM-top}
\beq{5.3.5b}
P_s(P_s +
 2\mu) = -\bar B_s/b_s,
\eeq
 $s \in S$, and the parameters $\bar B_s/b_s$  coincide within
blocks, i.e. $\bar B_s/b_s = \bar B_{s'}/b_{s'}$ for $s, s' \in
S_i$, $i =1,\dots,k$.

{\bf Finite-dimensional Lie algebras.}

Let $(A_{s s'})$ be  a Cartan matrix  for a  finite-dimensional
semisimple Lie  algebra $\cal G$. In this case all powers in
(\ref{5.2.20})  are  positive integers  which coincide with the
components of twice the  dual Weyl vector in the basis of simple
co-roots \cite{FS}, and  hence all functions $H_s$ are
polynomials, $s \in S$.

{\bf Conjecture 1.} {\em Let $(A_{s s'})$ be  a Cartan matrix for
a semisimple finite-dimensional Lie algebra $\cal G$. Then  the
solutions to eqs. (\ref{5.3.1})-(\ref{5.3.2b}) (if any) have a
polynomial structure:
 \beq{5.3.12} H_{s}(z) = 1 + \sum_{k =
                1}^{n_s} P_s^{(k)} z^k,
 \eeq
 where $P_s^{(k)}$ are constants, $k = 1,\ldots, n_s$; $n_s = b_s =
2 \sum_{s' \in S} A^{s s'} \in \N$ and $P_s^{(n_s)} \neq 0$,  $s
\in S$}.

In the extremal case ($\mu = + 0$), an analogue of this conjecture
was suggested previously in \cite{LMMP}. Conjecture 1 was verified
for the  ${\bf A_m}$ and ${\bf C_{m+1}}$ Lie algebras in
\cite{IMp2,IMp3}. Explicit expressions for polynomials
corresponding to the Lie algebras $C_2$ and $A_3$ were obtained in
\cite{GrIK} and \cite{GrIM}, respectively. Recently, a family of
black brane type solutions in the model  with multicomponent
anisotropic fluid were found in  \cite{I-2010}.

{\bf Hyperbolic KM algebras.} Let $(A_{s s'})$ be  a Cartan matrix
for the infinite-dimensional hyperbolic KM  algebra $\cal G$. In
this case, all powers in (\ref{5.2.20})  are  negative,  and hence
we have no chance to get a polynomial structure for $H_s$. Here we
are led to an open problem of seeking  solutions to the set of
``master'' equations (\ref{5.3.1})-(\ref{5.3.2a}). These solutions
determine special solutions to the Toda-chain equations
corresponding to the hyperbolic KM algebra $\cal G$.

{\bf Example 2. Black hole solutions for the KM algebras  $A_1
 \oplus A_1$, $A_2$ and  $H_2(q,q)$  \cite{IM-SIGMA}.}

Consider a 4-dimensional model governed by the action
\bear{5.4.1}
S =  \int_{M} d^{4}z \sqrt{|g|} \{ {R}[g] - \eps g^{MN}
\partial_{M} \varphi\partial_{N} \varphi -  \frac{1}{2}
e^{2\lambda \varphi} (F^1)^2 -  \frac{1}{2} e^{- 2\lambda \varphi}
(F^2)^2 \}.
\ear

Here $F^1$ and $F^2$ are 2-forms, $\varphi$ is a scalar field and
$\eps = \pm 1$.

We consider a black brane solution   defined on $\R_{*} \times
S^{2} \times \R $   with two electric branes   $s_1$ and $s_2$,
corresponding to the forms $F^1$ and $F^2$,   respectively, with
the sets $I_1 = I_2 = \{ 2 \}$.   Here $\R_{*}$ is a subset of
$\R$, $M_1 = S^{2}$,  $g^1 = d \Omega^2_{2}$  is the canonical
metric on $S^{2}$, $M_2 = \R$, $g^2 = - dt \otimes dt$ and $\eps_1
= \eps_2 = -1$.

The scalar products of the $U$-vectors are (we identify $U^i =
U^{s_i}$):

\bear{5.4.2} (U^1,U^1) = (U^2, U^2) = \frac{1}{2 } + \eps
\lambda^2  \neq 0,    \qquad (U^1,U^2) = \frac{1}{2 } - \eps
\lambda^2.  \ear

The matrix $A$ from (\ref{3.1.2.2}) is a generalized
non-degenerate Cartan matrix  if and only if
\beq{5.4.3}
\frac{2(U^1,U^2)}{(U^2, U^2)} = -q,
\eeq
 or, equivalently,
\beq{5.4.4}
\eps \lambda^2 = \frac{2+q}{2(2-q)},
\eeq
 where $q = 0,1,3,4, ...$. This takes place if $\eps = + 1$ for
$q = 0, 1$    and   $\eps = - 1$ for $q =  3, 4, 5, \dots$.

The first branch ($\eps = + 1$) corresponds to the
finite-dimensional Lie algebras $A_1 \oplus A_1$ ($q =0$), $A_2$
($q = 1$) and  the second one ($\eps = - 1$)  to the hyperbolic KM
algebras $H_2(q,q)$, $q = 3,4, ...$. In the hyperbolic case, the
scalar field $\varphi$ is a phantom.

The  black brane solution reads (see \ref{5.2.30})-(\ref{5.2.32}))

\bear{5.4.7}
g= (H_1 H_2)^{h} \biggl\{ \left(1 - \frac{2\mu}{R}
\right)^{-1} dR \otimes dR + R^2  d \Omega^2_{2}
\\ \nonumber
 - (H_1 H_2)^{- 2h}
\left(1 - \frac{2\mu}{R}\right) dt \otimes dt \biggr\},
\\  \label{5.4.8}
\exp(\varphi)= (H_1 / H_2)^{\eps \lambda h},
\\  \label{5.4.9}
F^s= \frac{Q_s}{R^2} H_s^{-2} (H_{\bar s})^{q} dt \wedge dR,
\ear
 $s = 1, 2$. Here $h = (2- q)/2$ and $\bar s = 2, 1 $ for $s = 1,
2$, respectively.

The moduli functions $H_s > 0$ obey  the equations (see
(\ref{5.3.1}))
\beq{5.4.10}
\frac{d}{dz} \left( \frac{(1 - 2\mu
z)}{H_s} \frac{d}{dz} H_s \right) = \frac{2 Q_s^2}{q-2} H_s^{-2}
(H_{\bar s})^{q},
\eeq
 with the boundary conditions  $H_{s}((2\mu)^{-1} -0) = H_{s0} \in
(0, + \infty)$,   $H_{s}(+ 0) = 1$, $s = 1, 2$. Here $\mu
 > 0$, $z = 1/R \in (0, (2\mu)^{-1})$. For $q = 0, 1$ the solutions
to eqs. (\ref{5.4.10}) with these boundary conditions  were given
in \cite{IMp1,IMp2,IMp3}. They are polynomials of degrees $1$ and
$2$ for $q=0$ and $q =1$, respectively. For $q = 3, 4, ...$ the
exact solutions to eqs. (\ref{5.4.10}) are yet unknown.

In the special case $Q_1^2 = Q_2^2$ the metric coincides with that
 of the Reissner-Nordstr\"om solution \cite{IM-SIGMA}.

{\bf Example 3: Black brane solution corresponding to the KM
algebra $HA_{2}^{(1)} = A_2^{++}$.}

Now we consider the $B_{15}$-model in  $15$-dimensional
pseudo-Euclidean space of signature $(-, +, ..., +)$ with the
forms $F^4$,..., $F^8$.

Here we deal with four electric branes   $s_1, s_2, s_3, s_4$
corresponding to the 6-form $F^6$.   The brane sets are: $I_1 = \{
1, 2, 3,  11, 12  \}$, $I_2 = \{ 4, 5, 6, 11, 12 \}$, $I_3 = \{ 7,
8, 9,  11, 12  \}$, $I_4 = \{ 1, 4, 10, 11, 12 \}$.

It may be verified that these sets obey the intersection rules
corresponding to  the hyperbolic KM algebra $HA_{2}^{(1)}$ with
the Cartan matrix
\bear{5a.H.A.2}
   A=\barr{cccc}
    2 & -1  & -1 &  0  \\
   -1 &  2  & -1 &  0  \\
   -1 & -1  &  2 & -1  \\
    0 &  0  & -1 &  2  \\
 \earr,
\ear
(see (\ref{3.1.2.3.sl}) with $I_{s_i} = I_i$).

Now we give a black brane solution  for the configuration of four
branes under consideration.

In what follows the relations  $\eps_s = +1$ and $h_s = 1/2$, $s
\in S$, are used.

 The metric (\ref{5.2.30}) reads
\bear{5a.3.ha.g}
g= (H_1 H_2 H_{3} H_4)^{5/13}
   \biggl\{ \left(1 -  \frac{2\mu}{R} \right)^{-1} dR \otimes dR + R^2  d \Omega^2_{2}
\\ \nonumber
 -  (H_1 H_2 H_{3} H_4)^{-1}
 \left[ \left(1 -  \frac{2\mu}{R} \right) dt^1 \otimes dt^1  + dt^2 \otimes dt^2 \right]
\\ \nonumber
 + (H_1 H_4)^{-1}  dx^1 \otimes dx^1
 + H_1^{-1} [ dx^2 \otimes dx^2   +  dx^3 \otimes dx^3 ]
\\ \nonumber
 + (H_2 H_4)^{-1}  dx^4 \otimes dx^4
 +  H_2^{-1} [ dx^5 \otimes dx^5
 +   dx^6 \otimes dx^6 ]
\\ \nonumber
 + H_3^{-1} [ dx^7 \otimes  dx^7
 +  dx^8 \otimes dx^8   +  dx^9 \otimes  dx^9 ]
\\ \nonumber
 + H_4^{-1}  dx^{10} \otimes dx^{10}
\biggr\}.
\ear

Here $t^1 = x^{11}$ and  $t^2 = x^{12}$ are time-like variables

The non-zero form field is
 \bear{5a.3.ha.f} F^6 =
 - Q_{1} R^{-2} H_1^{-2} H_2 H_3 dR \wedge dt^1 \wedge dt^2 \wedge dx^{1} \wedge dx^{2}\wedge dx^{3}
\\ \nonumber
 - Q_{2} R^{-2} H_1 H_2^{-2} H_3 dR \wedge dt^1 \wedge dt^2 \wedge dx^{4} \wedge dx^{5}\wedge dx^{6}
\\ \nonumber
 - Q_{3} R^{-2} H_1 H_2 H_3^{-2} H_4  dR \wedge dt^1 \wedge dt^2 \wedge dx^{7} \wedge dx^{8}\wedge dx^{9}
\\ \nonumber
 - Q_{4} R^{-2} H_3 H_4^{-2} dR \wedge dt^1 \wedge dt^2 \wedge dx^{1} \wedge dx^{4}\wedge dx^{10},
\ear
 where $Q_s \neq 0$, $s = 1, 2, 3, 4$.

 The scalar fields  read
\beq{5a.3.ha.p}
\varphi^{\alpha} =
 - \frac{1}{2} \lambda_{6 \alpha} \ln(H_1 H_2 H_{3} H_4),
\eeq
 $\alpha = 1,2,3,4$. Here   $H_s > 0$ obey the equations
 \beq{5a.3.1}
 \frac{d}{dz} \left( \frac{(1 - 2\mu z)}{H_s}
 \frac{d}{dz} H_s \right) = 2 Q_s^2
 \prod_{s' = 1}^4  H_{s'}^{- A_{s s'}}, \eeq
 with the boundary conditions
 $H_{s}((2\mu)^{-1} -0) = H_{s0} \in
(0, + \infty)$, and  $H_{s}(+ 0) = 1$, $s = 1, ..., 4$. Here
$\mu > 0$, $z = R^{-1} \in (0, (2\mu)^{-1})$  and   $(A_{s s'})$ is
the Cartan matrix (\ref{5a.H.A.2}) for the KM algebra
 $HA_2^{(1)}$.

{\bf Special 1-block solution.}
 This solution is valid when  a
special set of charges is considered:
\beq{4a.3.ha.Q}
 Q_s^2 = Q^2 |b_s|,
\eeq
 where $Q \neq 0$ and
\beq{4a.3.ha.b}
 b_s = 2 \sum_{s' =1}^{4} A^{ss'} = -12, -12, -14, - 6,
\eeq
for $s = 1, 2, 3, 4$, respectively. In this case the moduli
functions read \beq{5a.4.11}
 H_s  = H^{b_s}, \qquad  H =  1 + P/R,
\eeq

where $P(P + 2\mu) = 2Q^2$. These functions obey $H > 0$ for $R
\in [2 \mu, + \infty)$     if $P > - 2 \mu$ ($\mu > 0$). Due to
this inequality and the     relation  $P(P + 2 \mu) = Q^2$,
 we get
 \beq{5a.4.12}  P = -  \mu + \sqrt{\mu^2 + 2Q^2} > 0. \eeq

The Hawking temperature in this case is \cite{IM-top} $T_H = (1 +
P/2\mu)^{22}/(8 \pi  \mu)$. It diverges  as $\mu \to + 0$. This is
in agreement with the fact that the metric (\ref{5a.3.ha.g}) has a
singularity at $R = + 0$ if  $\mu = + 0$.

{\bf Example 4: Black brane solution corresponding to the
Lorentzian KM algebra  $P_{10}$.}

Now we consider another solution for  the $B_{15}$-model in
$15$-dimensional pseudo-Euclidean space of signature $(-, +, ...,
+)$   with the non-zero 6-form $F^6$.

Here we deal with ten electric branes   $s_1, ..., s_{10}$
corresponding to the 4-form $F^4$.   The brane sets are:
 $I_1 = \{
1, 4, 7, 11, 12  \}$, $I_2 = \{ 8, 9, 10, 11, 12 \}$, $I_3 = \{ 2,
5, 7, 11, 12  \}$, $I_4 = \{ 4, 6, 10, 11, 12 \}$, $I_5 = \{ 2, 3,
9, 11, 12  \}$, $I_6 = \{ 1, 2, 8, 11, 12 \}$, $I_7 = \{ 1, 3, 10,
11, 12 \}$, $I_8 = \{ 4, 5, 8, 11, 12 \}$, $I_9 = \{ 3, 6, 7, 11,
12  \}$, $I_{10} = \{ 5, 6, 9, 11, 12 \}$.

These sets obey  the intersection rules corresponding to the
Lorentzian KM algebra $P_{10}$ (we call it the Petersen algebra)
with the following Cartan matrix
\bear{5a.P.A.2}
A=\barr{cccccccccc}
    2 & -1  &  0 &  0 & -1 &  0 &  0 &  0  &  0 & -1 \\
   -1 &  2  & -1 &  0 &  0 &  0 &  0 &  0  & -1 &  0 \\
    0 & -1  &  2 & -1 &  0 &  0 & -1 &  0  &  0 &  0 \\
    0 &  0  & -1 &  2 & -1 & -1 &  0 &  0  &  0 &  0 \\
   -1 &  0  &  0 & -1 &  2 &  0 &  0 & -1  &  0 &  0 \\
    0 &  0  &  0 & -1 &  0 &  2 &  0 &  0  & -1 & -1 \\
    0 &  0  & -1 &  0 &  0 &  0 &  2 & -1  &  0 & -1 \\
    0 &  0  &  0 &  0 & -1 &  0 & -1 &  2  & -1 &  0 \\
    0 & -1  &  0 &  0 &  0 & -1 &  0 & -1  &  2 &  0 \\
   -1 &  0  &  0 &  0 &  0 & -1 & -1 &  0  &  0 &  2 .
\earr
 \ear

The Dynkin diagram for this Cartan matrix could be represented by
the Petersen graph  ``a star inside a pentagon''. $P_{10}$ is a
Lorentzian KM algebra. It is a subalgebra of $E_{10}$ \cite{HPS}.

Let us present a black brane solution  for the configuration of
$10$ electric branes under consideration.  The metric
(\ref{5.2.30}) reads

\bear{5p.3.p.g} g= \left(\prod_{s =1}^{10} H_s \right)^{5/13}
\biggl\{  \left(1 -  \frac{2\mu}{R} \right)^{-1} dR \otimes dR +
R^2  d \Omega^2_{2}
\\ \nonumber
- \left( \prod_{s =1}^{10} H_s \right)^{-1} \left[ \left(1 -
\frac{2\mu}{R} \right) dt^1 \otimes dt^1  + dt^2 \otimes dt^2
\right]
\\ \nonumber
+ (H_1 H_6 H_7)^{-1}  dx^1 \otimes dx^1
\\ \nonumber
+ (H_3 H_5 H_6)^{-1}  dx^2 \otimes dx^2 + (H_5 H_7 H_9)^{-1}  dx^3
\otimes dx^3
\\ \nonumber
+  (H_1 H_4 H_8)^{-1} dx^4 \otimes dx^4 +  (H_3 H_8 H_{10})^{-1}
dx^5 \otimes dx^5
\\ \nonumber
+  (H_4 H_9 H_{10})^{-1}  dx^6 \otimes dx^6 +  (H_1 H_3 H_9)^{-1}
dx^7 \otimes  dx^7
\\ \nonumber
+  (H_2 H_6 H_8)^{-1}   dx^8 \otimes dx^8 + (H_2 H_5 H_{10})^{-1}
dx^9 \otimes  dx^9
\\ \nonumber
+ (H_2 H_4 H_7)^{-1}   dx^{10} \otimes dx^{10} \biggr\}.
\ear

The form field  is

\bear{5p.3.p.f} F^6 =  - Q_{1} R^{-2} H_1^{-2} H_2 H_5 H_{10} dR
\wedge dt^1 \wedge dt^2  \wedge dx^{1} \wedge dx^{4}\wedge dx^{7}
\\ \nonumber
- Q_{2} R^{-2} H_1 H_2^{-2} H_3 H_9 dR \wedge dt^1 \wedge dt^2
\wedge dx^{8} \wedge dx^{9}\wedge dx^{10}
\\ \nonumber
- Q_{3} R^{-2} H_2 H_3^{-2} H_4 H_7 dR \wedge dt^1 \wedge dt^2
\wedge dx^{2} \wedge dx^{5}\wedge dx^{7}
\\ \nonumber
- Q_{4} R^{-2} H_3 H_4^{-2} H_5 H_6 dR \wedge dt^1 \wedge dt^2
\wedge dx^{4} \wedge dx^{6}\wedge dx^{10}
\\ \nonumber
- Q_{5} R^{-2} H_1 H_4 H_5^{-2} H_8 dR \wedge dt^1 \wedge dt^2
\wedge dx^{2} \wedge dx^{3}\wedge dx^{9}
\\ \nonumber
- Q_{6} R^{-2} H_4  H_6^{-2} H_9 H_{10} dR \wedge dt^1 \wedge dt^2
\wedge dx^{1} \wedge dx^{2}\wedge dx^{8}
\\ \nonumber
- Q_{7} R^{-2} H_3  H_7^{-2} H_8 H_{10} dR \wedge dt^1 \wedge dt^2
\wedge dx^{1} \wedge dx^{3}\wedge dx^{10}
\\ \nonumber
- Q_{8} R^{-2} H_5 H_7 H_8^{-2} H_9 dR \wedge dt^1 \wedge dt^2
\wedge dx^{4} \wedge dx^{5}\wedge dx^{8}
\\ \nonumber
- Q_{9} R^{-2} H_2 H_6 H_8 H_9^{-2} dR \wedge dt^1 \wedge dt^2
\wedge dx^{3} \wedge dx^{6}\wedge dx^{7}
\\ \nonumber
- Q_{10} R^{-2} H_1 H_6 H_7 H_{10}^{-2} dR \wedge dt^1 \wedge dt^2
\wedge dx^{5} \wedge dx^{6}\wedge dx^{9}, \ear
 where $Q_s \neq 0$, $s = 1, \dots, 10$.

The scalar fields  reads
\beq{5p.3.ha.p} \varphi^{\alpha} = -
\frac{1}{2} \lambda_{6 \alpha} \ln \left( \prod_{s =1}^{10} H_s
\right), \eeq $\alpha = 1,2,3,4$.
Here   $H_s > 0$ obey the
equations
 \beq{5p.3.1} \frac{d}{dz} \left( \frac{(1 - 2\mu
 z)}{H_s} \frac{d}{dz} H_s \right) = 2 Q_s^2  \prod_{s' = 1}^{10}
 H_{s'}^{- A_{s s'}},
 \eeq
 with the boundary conditions
$H_{s}((2\mu)^{-1} -0) = H_{s0} \in (0, + \infty)$, and  $H_{s}(+
0) = 1$, $s = 1, ..., 10$. Here $\mu > 0$, $z = R^{-1} \in (0,
(2\mu)^{-1})$,  and  $(A_{s s'})$ is the Cartan matrix
(\ref{5a.H.A.2})  for the KM algebra $P_{10}$.

{\bf Special 1-block solution.} Now we consider a special 1-block
solution. This solution is valid if a special set of charges is
considered:  $Q_s^2 = 2 Q^2$ ($Q \neq 0$) in agreement to
(\ref{4a.3.ha.Q}) and
 \beq{4p.3.ha.b} b_s = 2 \sum_{s' =1}^{10}
 A^{ss'} = -2, \eeq
 for $s = 1, ...,  10$.
In this case the  functions $H_s$ are
 \beq{5p.4.11} H_s  =
 H^{-2}, \qquad  H =  1 + P/R, \eeq
 where $P(P + 2\mu) = 2Q^2$. As
in the previous case, we get a well-defined   solution for $P = -
\mu + \sqrt{\mu^2 + 2Q^2} > 0$ and $\mu > 0$.

The Hawking temperature in this case has the following form: $T_H
 = (1 + P/2\mu)^{10}/(8 \pi  \mu)$. It is smaller than  that in the
previous example but it also diverges  as $\mu \to + 0$. It is in
agreement  with the  singularity of the metric (\ref{5p.3.p.g}) at
$R = + 0$ for   $\mu = + 0$.

 \section{Conclusions}

We have considered several  classes of exact solutions in
multidimensional gravity with a set of scalar fields and fields of
forms  related to non-singular (e.g., hyperbolic) KM algebras.

The solutions describe composite  electromagnetic branes defined
on warped products of  Ricci-flat, or sometimes Einstein, spaces
of arbitrary dimensions and signatures. The metrics are
block-diagonal, and all scale factors, scalar fields and fields of
forms depend on points of some  manifold $M_0$. The solutions
include those depending on harmonic functions, and
spherically-symmetric black-brane solutions. Our approach is based
on the  sigma-model representation  obtained in \cite{IMC} under
rather a general assumption on intersections of composite branes
(such that the stress-energy tensor has a diagonal structure).

We have also considered  a class of  black brane configurations
with   intersection rules \cite{IMJ} governed by an invertible
generalized Cartan matrix corresponding to a certain generalized
KM Lie algebra ${\cal G}$. The ``master'' equations for moduli
functions  have polynomial solutions in the finite-dimensional
case (according   to our conjecture \cite{IMp1,IMp2,IMp3}),  while
in the infinite-dimensional case we have only a special family of
the so-called block-orthogonal solutions  corresponding to
semi-simple  non-singular  KM algebras.   Certain examples  of
black brane solutions  are presented, corresponding to the
hyperbolic KM algebras: $H_2(q,q)$ ($q > 2$), $HA_2^{(1)} =
A_2^{++}$ and the Lorentzian KM algebra  $P_{10}$.

The last two solutions (which are new) may be analyzed from the
viewpoints (i) of post-Newtonian parameters $\beta$ and $\gamma$
corresponding to the 4-dimensional section of the metric and (ii)
of  thermodynamic properties of the black branes under
consideration. These and some other tasks may be a subject of a
separate publication.

\begin{center}
{\bf Acknowledgments}
\end{center}

This work was supported in part by grant NPK-MU (PFUR), Russian
Foundation for Basic Research  (Grant  Nr. 09-02-00677-a)  and by
the  FTsP ``Nauchnie i nauchno-pedagogicheskie kadry
innovatsionnoy Rossii''  for the years 2009-2013.

\end{document}